OpenStudy (anonymous):

SIMPLIFY THIS EXPRESSION!!!!!!! csc(-x)/1+tan^2(x) a) tan(x) b) -cos(x)tan(x) c) -cos(x)cot(x) d) sin(x)tan(x)

OpenStudy (anonymous):

tan^2 x+ 1 =sec^2 x

OpenStudy (anonymous):

csc (-x)= 1/six(-x)= -1/sin(x)

OpenStudy (anonymous):

sec (x) =1/cos (x)

OpenStudy (anonymous):

csc(-x)/sec^2(x) cos^2/sin (x)=-cos(x) cot(x)

OpenStudy (anonymous):

the answer is C